# Towards Physically-Realizable Adversarial Attacks in Embodied Vision Navigation

### v1

@@ IV-A Experimental Setup
TABLE I: Attack performance on multi-view object detection.
@@
Camouflage [2] 85.12(66.11$\uparrow$)
Adversarial Patch (Random Texture) 29.75(10.74$\uparrow$)
Adversarial Patch (2D Optimized Texture) 74.38(55.37$\uparrow$)
Ours Patch (Multi-view Optimized Texture) 89.26 (70.25$\uparrow$)
Ours Patch (Multi-view Optimized Texture & Opacity) 98.35 (79.34$\uparrow$)
TABLE II: Attack performance on the HM3D scene dataset for navigation.
@@
No Attack 19.01 63.03 0.04
Camouflage [2] 100.00(0.00$\downarrow$) 56.37(6.66$\downarrow$) 0.04(0.00$\uparrow$)
Ours Patch 60.00 (40.00$\downarrow$) 14.81 (48.22$\downarrow$) 1.68(1.64$\uparrow$)
Settings.  We use the modular-based navigation agent from [22] for ObjectNav, which includes a Mask R-CNN [44] for object instance segmentation, predicting 9 target categories from HM3D images. Our setup follows the 2022 Habitat ObjectNav Challenge [45]. The adversarial patch is $512\times 512$, initialized with Gaussian noise, and randomly placed on the target. Mitsuba 3 [46] serves as the renderer, with planar area lights or “constant” scene lighting at intensity 40, and a camera resolution of $512\times 512$. The detection confidence threshold is 0.5. We apply PGD (L2/Linf) attacks with a maximum perturbation of 5 units over 100 steps, with a learning rate of 1.
Datasets.  We conduct adversarial attack experiments on ObjectNav and the HM3D dataset [47] using the Habitat simulator [48, 20]. The TV monitor is chosen as the attack target, with one scene from the validation dataset used for white-box optimization, targeting a $512\times 512$ transparent patch.
Evaluations.  We evaluate attack performance using three metrics from [22]: Success Rate (SR), Success weighted by Path Length (SPL), and Distance to Goal (DTS). SR measures the agent’s ability to successfully locate the target object, while SPL considers both the success rate and the efficiency of the path taken. DTS quantifies the remaining distance between the agent and the goal at the end of each episode. Additionally, we measure the Attack Success Rate (ASR) for multi-view object detection in the scene.
IV-B Attack Performance
@@
Figure 4: Visualization of the attack performance of our adversarial patch compared to 3D camouflage in the navigation task. The first row shows the adversarial scenario with 3D camouflage, where the agent was still able to successfully navigate and locate the target. In contrast, in the scenario with our adversarial patch, the agent failed to detect the target even after 499 steps, resulting in navigation failure. This highlights the superior attack performance of our method compared to the baseline 3D camouflage approach.
Figure 5: Visualization of adversarial patches under different ablation conditions. We display six viewpoints from left to right. From top to bottom, ”Random Noise” applies random textures without opacity, making the patch more visible in the scene. ”Optimized Texture” refines the texture through multi-view optimization, significantly altering the texture. ”Optimized Opacity” further optimizes the opacity, ”No Attack” represents the original clean scene. It is evident that after applying opacity optimization, the Natural of the patch improves significantly, aligning with the goals of physical adversarial attacks: effective performance, easy implementation, and reduced visibility to the human eye.
Table I presents the Attack Success Rate (ASR) for multi-view target detection in a navigation scenario, where 121 viewpoints surrounding the target were sampled using our object-aware sampling strategy. We compared our method with previous physically realizable approaches, specifically the camouflage adversarial patch method. The camouflage approach is based on [2], where we focused on the Spatially Contextual Perturbations module and incorporated multi-view optimization, as the Temporal Attention Stimulus module is not applicable to navigation tasks. The adversarial patch [3, 4, 5, 6, 7, 8, 9, 10, 11] was compared using random textures and 2D-optimized textures without differentiable rendering and multi-view optimization. Since direct placement of the patch on the target can be too conspicuous, we set the opacity to 0.8. Our experiments show that the proposed multi-view sampling strategy increased ASR by up to 70.25%. With the addition of opacity optimization, ASR further improved, reaching up to 79.24%, surpassing the baseline performance.
Table II shows the attack performance of our adversarial patch, which underwent multi-view optimization and opacity optimization, in the navigation task. Since camouflage [2] is the only method targeting embodied navigation, we limited our comparison to it. Our method reduced the navigation success rate by approximately 40% in the surrogate scene. The visualization and qualitative analysis in Fig. 4 further illustrate the patch’s effectiveness, where the agent failed to detect the target even after 499 steps, resulting in navigation failure. This demonstrates superior attack performance compared to the baseline 3D camouflage method.
@@ IV-C Ablation Study
Effect of multi-view optimization based on object-aware sampling. The experimental results are shown in Table III. As observed, when neither multi-view texture optimization nor opacity optimization is applied, the ASR is the lowest at 28.93%. When only single-view texture optimization is applied, there is an improvement, and multi-view texture optimization leads to even higher performance. When both texture and opacity are optimized, the ASR reaches its highest value of 99.17%, validating the effectiveness of our multi-view optimization strategy. The visualizations are presented in Fig. 5, where we display six viewpoints from left to right. From top to bottom, we show the random texture, optimized texture, optimized opacity, and the original scene, respectively. It is evident that after applying opacity optimization, the visual quality of the patch improves significantly, aligning with the goals of physical adversarial attacks: easy to implement, effective performance, and less visible to the human eye.
Effect of opacity value. In addition, we studied the effect of different opacity levels in Table IV. We applied varying levels of opacity to both random and optimized textures. At an opacity of 0.2, the patch is almost completely transparent and nearly invisible to the human eye, while at 0.8, it is nearly fully opaque. The higher the opacity, the higher the ASR, but the patch becomes more noticeable. Therefore, we selected an initial opacity of 0.6 for both texture and opacity optimization in our quantitative experiments, as it provides a balance between attack effectiveness and visibility. Finally, in Fig. 6, we present the rendered effects of the adversarial patch with different opacity values under various renderers.
@@
- 28.93
1 73.55(44.62$\uparrow$)
121 89.26(60.33$\uparrow$)
121 99.17(70.24$\uparrow$)
@@
V Conclusion
In this paper, we address the challenge of adversarial robustness in embodied navigation tasks and the physical feasibility of attack methods. We propose a practical attack method for embodied navigation by attaching adversarial patches with learnable textures and opacity to objects. Specifically, to ensure effectiveness across varying viewpoints, we employ a multi-view optimization strategy based on object-aware sampling, which uses feedback from the navigation model to optimize the patch’s texture. To make the patch inconspicuous to human observers, we introduce a two-stage opacity optimization mechanism, where opacity is refined after texture optimization. Experimental results show our adversarial patches reduce navigation success rates by about 40%, outperforming previous methods in practicality, effectiveness, and naturalness.

### v5

@@ IV Experiments
TABLE I: Attack performance against multiple targets in HM3D for navigation
[…]
tv monitor No Attack 100.00 63.03 0.04
[…]
plant No Attack 93.75 26.60 0.24
[…]
bed No Attack 94.11 55.83 0.16
[…]
chair No Attack 100.00 65.18 0.22
[…]
sofa No Attack 100.00 57.28 0.36
[…]
toilet No Attack 81.25 43.32 1.32
[…]
Average No Attack 94.85 51.87 0.39
[…]
@@ IV-A Experimental Setup
Settings. We use the modular-based navigation agent from [38] for ObjectNav, which incorporates a Mask R-CNN [39] for object instance segmentation. Our navigation experiment setup follows the 2022 Habitat ObjectNav Challenge [40]. For the adversarial attack, we conduct experiments in ObjectNav using the HM3D dataset [41] within the Habitat simulator [42, 43]. We perform white-box optimization using scenes from the validation dataset, with attack targets being several categories listed in Table I.
A $512\times 512$ adversarial patch with four RGBA channels is used, initialized with Gaussian noise and manually positioned on the target object. Mitsuba 3 [44] serves as the renderer, using planar area lights or “constant” scene lighting with an intensity parameter of 40., and a camera resolution of $512\times 512$. During texture and opacity optimization, the detection confidence threshold is set to 0.5. We apply PGD attacks over 100 iterations with a step size of 1.
For each target object category, we randomly select one scene from the HM3D validation set that contains the object and run 100 test episodes with random initial positions. The high success rates in the no-attack condition reflect the strong performance of the modular navigation pipeline under these controlled evaluation settings.
Evaluations. We evaluate attack performance using three metrics from [38]: Success Rate (SR), Success weighted by Path Length (SPL), and Distance to Goal (DTS). SR measures the agent’s ability to successfully locate the target object, while SPL considers both the success rate and the efficiency of the path taken. DTS quantifies the remaining distance between the agent and the goal at the end of each episode. Additionally, we measure the Attack Success Rate (ASR), which is defined as the proportion of viewpoints that were successfully attacked (i.e., viewpoints where detection fails) among all sampled viewpoints.
IV-B Attack Performance
TABLE II: Attack performance on multi-view object detection.
@@ IV-B Attack Performance
[…]
Table I presents the attack performance of our adversarial patch with multi-view and opacity optimization. As camouflage [8] is the only physically realizable attack tailored for embodied navigation, we use it as our baseline. Our method achieves an average SR reduction of 22.39% across various targets, along with notable decreases in SPL and increases in DTS, indicating less efficient and more difficult navigation. Figure 4 visualizes the stronger disruption effect of our approach, while Figure 5 highlights its superior visual naturalness.
@@
Figure 4: Visualization of the attack performance of our adversarial patch compared to 3D camouflage in the navigation task. The first row shows the adversarial scenario with 3D camouflage, where the agent was still able to successfully navigate and locate the target. In contrast, in the scenario with our adversarial patch, the agent failed to detect the target even after 499 steps, resulting in navigation failure.
Table II reports the Attack Success Rate (ASR) for multi-view target detection. We applied our object-aware sampling strategy around a television monitor, generating 120 viewpoints—20 for optimization and 100 unseen for testing. ASR is defined as the proportion of failed detections among these 100 views (confidence threshold 0.5). We compared our method with prior physically realizable attacks, including camouflage [8], 2D-optimized patches without differentiable rendering [16, 17, 18, 19, 20, 21], and randomly initialized textures. Our method achieved 89.26% ASR, which further increased to 98.35% with opacity optimization, clearly outperforming all baselines.
Figure 5: Comparison of naturalness. From left to right, the images show the previous camouflage method, our adversarial patch, and the clean scene. Our adversarial patch exhibits greater visual naturalness.
@@ IV-C Ablation Study
- 19.01
[…]
Effect of multi-view optimization based on object-aware sampling. The experimental results are shown in Table III. As observed, when neither multi-view texture optimization nor opacity optimization is applied, the ASR is the lowest at 19.01%. When only single-view texture optimization is applied, there is an improvement, and multi-view texture optimization leads to even higher performance. When both texture and opacity are optimized, the ASR reaches its highest value of 98.35%, validating the effectiveness of our multi-view optimization strategy.
Figure 6: Visualization of the effect of opacity optimization. It is evident that after applying opacity optimization, the naturalness of the patch improves significantly.
Effect of opacity optimization. As shown in Figure 6, opacity optimization significantly improves the patch’s visual naturalness. We further examined different initial opacity levels (Table IV) on both random and optimized textures. Lower opacity (e.g., 0.2) renders the patch nearly invisible, while higher values (e.g., 0.8) improve ASR but reduce stealth. This positive correlation mainly holds for optimized textures; random textures lack alignment with model vulnerabilities, making opacity less effective. We selected an initial opacity of 0.6 to balance attack strength and visibility, and then refined the mask accordingly. Figure 7 shows the patch’s appearance under different renderers.
Physically, the patch can be printed on transparent paper with a laser printer and directly affixed to target objects, avoiding any modification to the scene geometry. Its adjustable transparency further enhances stealth in real-world navigation scenarios.
@@
V Conclusions
In this paper, we propose a practical attack on embodied navigation by applying adversarial patches with learnable textures and opacity to objects, reducing success rates by 22.39% on average. While the method shows strong performance in simulation, future work is needed to improve patch printing techniques and evaluate human perceptibility in real-world settings. Despite the lack of physical experiments, our approach outperforms prior methods—including first-person perturbations, full 3D texture attacks, and less effective multi-view patches—in physical feasibility, multi-view robustness, and visual naturalness. Moreover, Virtual 3D environments continue to play a key role in advancing embodied navigation research.
Acknowledgment
This work was supported by the Beijing Natural Science Foundation (Grant No. F2024203115), the National Natural Science Foundation of China (Grant No. 62173045), the National Natural Science Foundation of China (Grant No. 62403491), the China Postdoctoral Science Foundation (Grant No. 2024M750255), the “Double First-Class” Interdisciplinary Team Project of Beijing University of Posts and Telecommunications (Grant No. 2023SYLTD02), and the National Key R&D Program of China (Grant No. 2024YFC3015604).